find the area enclosed by the curve r^2=4cos theta


1
Expert's answer
2022-06-06T14:43:51-0400

Let's graph the curve "r^2=4cos \theta."



The graph of the curve is symmetrical to the origin and the area consists of 4 equal parts.

So we can find the area of one part ("\theta" ranges from 0 to "\frac{\pi}{2}") and multiply it by 4:

"A=4\cdot\frac{1}{2}∫_{0}^{\frac{\pi}{2}}[r(\theta)]^2dθ="

"=2∫_{0}^{\frac{\pi}{2}}4cos \theta dθ=8∫_{0}^{\frac{\pi}{2}}cos \theta dθ="

"=8 sin \theta |_0^{\frac{\pi}{2}}=8(sin\frac{\pi}{2}-sin0)=8(1-0)=8."


Answer: 8 square units.